# Cosmological Simulations of Stellar Halos with Gaia Sausage–Enceladus Analogues: Two Sausages, One Bun?

### v1

@@ 2.3 Recovering the Merger History
To study the assembly history of each MW analogue’s ex situ stellar population, we build a pipeline that appropriately identifies the progenitor for each star in its halo. The key points of this procedure are summarized in this subsection, with further details provided in Appendix B.
@@
The stars considered in this study are those which are either: (i) bound directly to the MW analogue at $z=0$, or (ii) bound to a non-cosmological, primarily-gaseous overdensity identified by the clustering algorithm as an independent gravitationally-bound structure within the MW analogue. This latter case is noted in the TNG catalog under the SubhaloFlag field, as described on the data release page. [^4] https://tng-project.org/data/docs/background We consider all such overdensities within the MW analogue’s $R_{\mathrm{dyn}}$ and add their stars to the sample.
@@
For a star in the MW analogue, we locate its unique particle ID in every snapshot of the simulation and record the halos to which it is bound. [^5] We exclude snapshots before the formation of the star particle, as well as snapshots in which Subfind does not find the particle to be gravitationally bound to any halo. With this information, we associate stars to the mergers which contributed them, accounting for both mergers which are completely disrupted and those which survive to $z=0$ (see Appendix B for more information). The masses $M_{\mathrm{dyn}}$ and $M_{\star}$ are recorded for each merger at every snapshot. Each merger is assigned an infall time, which is the last time the halo crosses its MW analogue’s $R_{\mathrm{dyn}}$.
@@
Stars are tagged as originating from the merger to which they are bound for the greatest number of snapshots, with preference given to the one with the largest peak mass in cases of ambiguity. [^6] Ambiguous cases are rare—see Appendix B. For a star to be considered ex situ, the following criteria must be met:
@@
  • in the first snapshot after the star formed, it must be bound to a halo which is not the MW analogue itself, but which is flagged as cosmological in origin according to the SubhaloFlag.
@@
These criteria prioritize the purity of the ex situ stellar sample. Another potential definition of ex situ material is a simple cut on the distance from the host at the time the star particle formed, but this is more challenging to apply to the TNG hosts whose stellar disks (and therefore in situ populations) are often quite extended (Pillepich et al., 2023). As a result, typical scales used for birth distance selections may not be easily applicable. The second ex situ criterion used in this work sidesteps this complication but is in the same spirit, as the ex situ stars must not be bound to the MW analogue at the time of their formation.
@@
Figure 1: (Left:) The composition of the ex situ stellar halo for a particular MW analogue (ID 543729) as a function of galactocentric distance $r$. The five largest galaxies contributing to the ex situ stellar population are shown as individual colored bands, sorted by the contribution within 40 kpc. The galaxy which contributes the most ex situ stellar mass is colored in blue, the second in purple, the third in magenta, the fourth in orange, and the fifth in gold. Stars contributed by all other mergers are collected into the light gray band at the top. This MW analogue’s ex situ population is primarily from one large merger with infall $M_{\mathrm{dyn}}\sim 8\times 10^{10}~{}\mathrm{M}_{\odot}$. All other mergers contribute fewer stars to the accreted population. The horizontal line indicates 50%, the threshold used throughout the work to denote significant mergers. (Right:) The same as the left panel, but for ID 516760. This halo hosts a pair of mergers (blue and purple) that, taken together, comprise more than 50% of the ex situ stellar material within 40 kpc.
[…]
*4
3 Stellar Halos in TNG
@@
A population of 98 MW-like galaxies passes the cuts established in §2.2. These galaxies have total masses of [^7] Logarithms are taken to be base ten throughout this work. $\log(M_{\mathrm{dyn}}/\mathrm{M}_{\odot})=12.05^{+0.17}_{-0.11}$, corresponding to $R_{\mathrm{dyn}}=220^{+30}_{-18}$ kpc (16, 50, 84th percentile). Generally, these galaxies are central to their FoF groups; however, four analogues are part of cluster-scale groups, though they still satisfy the isolation criteria.
@@
All MW analogues are relatively devoid of large satellites, with only 1 (11) analogue hosting a satellite of $M_{\mathrm{dyn}}>10^{10}~{}\mathrm{M}_{\odot}$ within 50 (100) kpc. This is in relative agreement with both theoretical work—e.g. Boylan-Kolchin et al. (2010); Busha et al. (2011); Bose et al. (2020); Buch et al. (2024)—and observations—e.g. Liu et al. (2011); Tollerud et al. (2011); Mao et al. (2024)—which find about ten percent of MW halos to host such bright dwarfs (though varying mass and distance criteria shift this number).
@@
In addition to searching for analogues with LMC partners, one can also classify those that experienced a GSE-like merger(s). In practice, this is more challenging, as it requires tracking the origin of the disrupted material within the TNG halos, rather than simply querying surviving satellites. The merger identification algorithm developed in this work (§2.3) enables one to build such histories. This section describes the accreted stellar material in these MW analogues, focusing on identifying those with RA debris expected from GSE-type events. Note that, henceforth, “stellar halo” is used to mean “accreted stellar population,” excluding in situ stars and with no restriction on kinematics.
3.1 Composition of the Stellar Halos
@@
Observations of the MW’s stellar halo indicate that the primary source of accreted stellar material in the MW is a small number of dwarf galaxies, with the largest identified component, the GSE, comprising as much as 50% of the accreted stars within 40 kpc [^8] This is true as a whole, but not necessarily differentially. Specifically, at small radii ($\lesssim 5$ kpc), stars from other accreted structures, such as Heracles (Horta et al., 2021), contribute more than the GSE; similarly, at larger radii ($\gtrsim 35$ kpc), the contribution from the GSE becomes small with respect to the Sagittarius dSph (Naidu et al., 2020). of the Galactic center (Deason et al., 2015, 2018; Necib et al., 2019a; Naidu et al., 2020). Previous theoretical studies—e.g. Deason et al. (2016); D’Souza & Bell (2018); Elias et al. (2020); Santistevan et al. (2020); Fattahi et al. (2020)—indicate that this may be a generic prediction of structure formation in the CDM paradigm, but the large volume and high resolution of the TNG simulation allow for a new look at the subject with unprecedented statistical power.
@@
The panels of Figure 1 show the ex situ stellar halo composition for two example MW analogues in the TNG set, as built from the tagging procedure of §2.3. Each of the five largest mergers that contribute to the halo is indicated by a different color, ranging from the blue merger, which contributes the most stellar mass, to the gold merger, which contributes the fifth most. The contribution from all remaining mergers is shown in gray. A dashed black line indicates 50%, the threshold used later to select significant mergers from the stellar halo. The halo shown in the left panel is overwhelmingly built of the debris of a single merger at radii $\lesssim 40$ kpc. The stellar halo shown in the right panel, however, has two quite significant mergers which together comprise half of the accreted stellar population at all radii $\lesssim 15$ kpc.
@@
For most stellar halos in the TNG sample (63%), over half of the accreted stars within 40 kpc come from a single merger. For most other halos (33%), summing contributions from two mergers is sufficient to reach the 50% threshold, and for a minority (4%) of the halos it takes three mergers to reach this threshold, with no halos requiring four or more mergers. These statistics are summarized in the left panel of Figure 2, which shows the fraction of MWs for which 50% of the stellar halo is comprised of one, two, and three mergers. It is therefore not rare for a MW analogue’s halo to be comprised predominantly of a single merger, but it is also not rare for two mergers to together comprise the majority of its ex situ population. In this work, we consider GSE analogues from both of these cases.
@@
Figure 2: (Left:) The number of mergers required to reach the threshold of 50% of ex situ stars within 40 kpc of the galactic center. Approximately 60% of the MW analogues have a single merger which contributes over half of the ex situ stars within 40 kpc, and an additional $\sim 30\%$ reach this threshold with two mergers. It is therefore not unreasonable to expect that an ex situ halo could be predominantly comprised of two mergers, though typically there is a single large merger which comprises the majority of the inner halo. (Right:) The velocity anisotropy $\beta$ of ex situ stellar debris is shown for all large mergers (those which contribute over 50% of their host’s ex situ stellar population within 40 kpc) in red. Those which are considered radially-anisotropic (RA) debris lie in the shaded large-$\beta$ region. Each pair of mergers which does sum to $f>50\%$ is shown in purple. There are occasionally multiple choices of pairs of mergers in the same halo that satisfy the RA debris criteria (i.e., that fall in the shaded region). There are 21 single-halo RA mergers, and there are 16 RA pairs of mergers selected from 11 MW analogues (5 MW analogues have multiple choices of RA pair).
3.2 GSE Analogue Selection
@@
Of the mergers identified in TNG, we define a GSE-like subsample of merger debris which
@@
An individual merger which satisfies both criteria is deemed a “RA merger.” In the halos where no single merger yields $f>0.5$ (regardless of its anisotropy), pairs of mergers are considered. In these cases, any two mergers whose stars together satisfy $f>0.5$ and $\beta>0.5$ are deemed a “RA pair.” Stars which are from either a RA merger or a RA pair are referred to as “RA debris.”
These criteria are generally influenced by considerations of the GSE debris in our Galaxy, as the GSE comprises the majority of ex situ material in the inner halo and has a high velocity anisotropy. Criterion (i) ensures that the GSE analogue contributes significantly to the inner halo, with no other large contributions, as inferred observationally. Criterion (ii) ensures the debris qualitatively resembles the GSE kinematically, though this criterion is especially permissive.
Of the 98 MW analogues, 21 have a single RA merger and 11 have a RA pair. [^9] We have verified that the RA pairs are not accreted together as part of a larger group structure. The choice of RA pair is not necessarily unique: in five halos, there are two combinations which each satisfy the requirements of a RA pair. Each combination is taken independently, though there is consistently one merger shared between the two pairs. All pairs will be specially marked in the figures, with icons denoting the MW analogues to which they belong.
@@
Notably, individual mergers tend not to have as extreme an anisotropy as the MW stellar halo; the GSE debris in the MW has $\beta\sim 0.8$ (Belokurov et al., 2018; Myeong et al., 2018; Lancaster et al., 2019; Iorio & Belokurov, 2021), which is at the far upper end of what is seen in TNG. Large single mergers have $\beta>0.8$ approximately 5% of the time, while pairs of mergers only have a $\beta$ this large 2.5% of the time. This implies the high anisotropy seen in the MW is quite rare.
Considering two mergers together as one set of debris tends to reduce the measured anisotropy, as the mergers do not necessarily have any correlation in velocity space; this is reflected in the shift toward isotropy ($\beta=0$) in the right panel of Figure 2. Despite this, there are still many pairs of mergers for which the contributed stars satisfy both the high-$\beta$ and high-$f$ requirements. This suggests that the presence of such an anisotropic structure cannot immediately preclude the possibility of a multi-component model. The rest of this work characterizes and aims to distinguish individual GSE-like RA mergers from RA pairs.
@@
3.3 Progenitors of GSE-like Debris
Various properties of GSE analogues are shown in Table 2, for both the RA mergers (top section) and RA pairs (bottom section). The columns include
  • MW ID, the Subfind ID of the MW analogue to which the GSE analogue belongs;
  • $f$, the fractional contribution of ex situ stars in the inner 40 kpc belonging to the GSE analogue;
@@ 3.3 Progenitors of GSE-like Debris
Figure 3: The infall times $t_{\mathrm{infall}}$ and masses $M_{\mathrm{dyn}}$ of all mergers across the TNG MW analogues. The RA mergers passing the selection criteria established in §3.2 are shown in red, while RA pairs are shown in purple (for the larger-$f$ merger) and pink (for the smaller-$f$ merger), with marker shapes indicating the MW analogue to which they belong. Other (non-RA) mergers are shown in black, with the opacity set by $f$. The corresponding marginal distributions are shown above and to the right of the main panel. Generally, RA mergers are more recently accreted and larger in mass than those comprising RA pairs. For the RA pairs, the larger-$f$ mergers tend to be more recently accreted and to have slightly larger mass.
@@
Figure 4: The distribution of spherical velocities $v_{r}$ and $v_{\phi}$ for RA merger debris in four example MW analogues. The grayscale histogram shows the full distribution of RA merger debris. The top row shows these distributions for single RA mergers, with 68% of the stars enclosed within the red contour. The bottom row shows the distributions for RA pairs, with 68% of stars from the high-$f$ (low-$f$) merger enclosed within the purple (pink) contour. In each bottom panel, the fractions $f$ are provided in the lower-left corner and an icon unique to the MW analogue is provided in the lower-right corner, consistent with Figure 3. The examples chosen here highlight the diversity in kinematic features observed in MW’s with RA debris. Versions of this figure for each choice of RA debris are included in Appendix C.
@@
Figure 5: Total energy shown against the $z$-component of angular momentum for the same MW analogues as in Figure 4, with identical formatting, except as noted. Each axis has been normalized to an energy and angular momentum scale characteristic of the RA debris to emphasize the shapes of these distributions. Specifically, $\bar{E}$ is the median energy of RA debris, while $\Delta E$ is the difference between the 2nd and 98th percentile. Similarly, $\Delta L_{z}$ is the difference between the 2nd and 98th percentile values of RA debris’ $L_{z}$. As in the $v_{r}-v_{\phi}$ plane, these distributions exhibit varying degrees of complexity irregardless of whether they correspond to single or pairs of mergers. Versions of this figure for each choice of RA debris are included in Appendix C.
@@
Across the sample, there are only three cases in which an infalling GSE-like galaxy survives to $z=0$. This happens once for a single RA merger (in ID 515296) and twice for the larger-$f$ partner of an RA pair (in MWs 392277 and 394621). In all cases where the GSE-like debris is built up predominantly from a single merger, that event is the most recent major merger (i.e., no subsequent merger has larger $M_{\mathrm{dyn}}/\mathrm{MW}~{}M_{\mathrm{dyn}}$). However, there are five MWs which accrete an RA pair and subsequently have a merger which is more significant than either in the pair. This happens in IDs 392277, 394621, 446665, 459557, and 535050.
@@
Further elements of the table are summarized in Figure 3, which shows the distribution of infall times, $t_{\mathrm{infall}}$, and total masses at infall, $M_{\mathrm{dyn}}$, for RA mergers (in red), RA pairs (in purple and pink, with the larger-$f$ merger in purple), and all other mergers (in black, with opacity set by the fraction $f$) across the sample. The marginal distributions of these parameters for the RA debris are shown as normalized histograms on each axis.

### v2

@@ II.3 Recovering the Merger History
To study the assembly history of each MW analog’s ex situ stellar population, we build a pipeline that appropriately identifies the progenitor for each star in its halo. The key points of this procedure are summarized in this subsection, with further details provided in B.
@@
The stars considered in this study are those that are either (i) bound directly to the MW analog at $z=0$, or (ii) bound to a noncosmological, primarily gaseous overdensity identified by the clustering algorithm as an independent gravitationally bound structure within the MW analog. This latter case is noted in the TNG catalog under the SubhaloFlag field, as described on the data release page. [^4] https://tng-project.org/data/docs/background We consider all such overdensities within the MW analog’s $R_{\mathrm{dyn}}$ and add their stars to the sample.
@@
For a star in the MW analog, we locate its unique particle ID in every snapshot of the simulation and record the halos to which it is bound. [^5] We exclude snapshots before the formation of the star particle, as well as snapshots in which Subfind does not find the particle to be gravitationally bound to any halo. With this information, we associate stars with the mergers that contributed them, accounting for both mergers that are completely disrupted and those that survive to $z=0$ (see B for more information). The masses $M_{\mathrm{dyn}}$ and $M_{\star}$ are recorded for each merger at every snapshot. Each merger is assigned an infall time, which is the last time the halo crosses its MW analog’s $R_{\mathrm{dyn}}$.
@@
Stars are tagged as originating from the merger to which they are bound for the greatest number of snapshots, with preference given to the one with the largest peak mass in cases of ambiguity. [^6] Ambiguous cases are rare—see B. For a star to be considered ex situ, the following criteria must be met:
  1. 1.
@@
    in the first snapshot after the star formed, it must be bound to a halo that is not the MW analog itself, but that is flagged as cosmological in origin according to the SubhaloFlag.
@@
These criteria prioritize the purity of the ex situ stellar sample. Another potential definition of ex situ material is a simple cut on the distance from the host at the time the star particle formed, but this is more challenging to apply to the TNG hosts whose stellar disks (and therefore in situ populations) are often quite extended (Pillepich et al., 2023). As a result, typical scales used for birth distance selections may not be easily applicable. The second ex situ criterion used in this work sidesteps this complication but is in the same spirit, as the ex situ stars must not be bound to the MW analog at the time of their formation.
@@
Figure 1: (Left:) The composition of the ex situ stellar halo for a particular MW analog (ID 543729) as a function of galactocentric distance $r$. The five largest galaxies contributing to the ex situ stellar population are shown as individual colored bands, sorted by the contribution within 40 kpc. The galaxy that contributes the most ex situ stellar mass is colored in blue, the second in purple, the third in magenta, the fourth in orange, and the fifth in gold. Stars contributed by all other mergers are collected into the light gray band at the top. This MW analog’s ex situ population is primarily from one large merger with infall $M_{\mathrm{dyn}}\sim 8\times 10^{10}~\mathrm{M}_{\odot}$. All other mergers contribute fewer stars to the accreted population. The horizontal line indicates 50%, the threshold used throughout the work to denote significant mergers. (Right:) The same as the left panel but for ID 516760. This halo hosts a pair of mergers (blue and purple) that, taken together, comprise more than 50% of the ex situ stellar material within 40 kpc.
III Stellar Halos in TNG
@@
A population of 98 MW-like galaxies passes the cuts established in Section II.2. These galaxies have total masses of [^7] Logarithms are taken to be base ten throughout this work. $\log(M_{\mathrm{dyn}}/\mathrm{M}_{\odot})=12.05^{+0.17}_{-0.11}$, corresponding to $R_{\mathrm{dyn}}=220^{+30}_{-18}$ kpc (16th, 50th, and 84th percentiles). Generally, these galaxies are central to their FoF groups; however, four analogs are part of cluster-scale groups, though they still satisfy the isolation criteria.
@@
All MW analogs are relatively devoid of large satellites, with only 1 (11) analog hosting a satellite of $M_{\mathrm{dyn}}>10^{10}~\mathrm{M}_{\odot}$ within 50 (100) kpc. This is in relative agreement with both theoretical work—e.g. Boylan-Kolchin et al. (2010); Busha et al. (2011); Bose et al. (2020); Buch et al. (2024)—and observations—e.g. Liu et al. (2011); Tollerud et al. (2011); Mao et al. (2024)—which find that about 10% of MW halos host such bright dwarfs (though varying mass and distance criteria shift this number).
@@
In addition to searching for analogs with LMC partners, one can also classify those that experienced a GSE-like merger(s). In practice, this is more challenging, as it requires tracking the origin of the disrupted material within the TNG halos, rather than simply querying surviving satellites. The merger identification algorithm developed in this work (Section II.3) enables one to build such histories. This section describes the accreted stellar material in these MW analogs, focusing on identifying those with RA debris expected from GSE-type events. Note that, henceforth, “stellar halo” is used to mean “accreted stellar population,” excluding in situ stars and with no restriction on kinematics.
III.1 Composition of the Stellar Halos
@@
Observations of the MW’s stellar halo indicate that the primary source of accreted stellar material in the MW is a small number of dwarf galaxies, with the largest identified component, the GSE, comprising as much as 50% of the accreted stars within 40 kpc [^8] This is true as a whole but not necessarily differentially. Specifically, at small radii ($\lesssim 5$ kpc), stars from other accreted structures, such as Heracles (Horta et al., 2021), contribute more than the GSE; similarly, at larger radii ($\gtrsim 35$ kpc), the contribution from the GSE becomes small with respect to the Sagittarius dSph (Naidu et al., 2020). of the Galactic center (Deason et al., 2015, 2018; Necib et al., 2019a; Naidu et al., 2020). Previous theoretical studies—e.g. Deason et al. (2016); D’Souza and Bell (2018); Elias et al. (2020); Fattahi et al. (2020); Santistevan et al. (2020)—indicate that this may be a generic prediction of structure formation in the CDM paradigm, but the large volume and high resolution of the TNG simulation allow for a new look at the subject with unprecedented statistical power.
@@
The panels of Figure 1 show the ex situ stellar halo composition for two example MW analogs in the TNG set, as built from the tagging procedure of Section II.3. Each of the five largest mergers that contribute to the halo is indicated by a different color, ranging from the blue merger, which contributes the most stellar mass, to the gold merger, which contributes the fifth most. The contribution from all remaining mergers is shown in gray. A dashed black line indicates 50%, the threshold used later to select significant mergers from the stellar halo. The halo shown in the left panel is overwhelmingly built of the debris of a single merger at radii $\lesssim 40$ kpc. The stellar halo shown in the right panel, however, has two quite significant mergers that together comprise half of the accreted stellar population at all radii $\lesssim 15$ kpc.
@@
For most stellar halos in the TNG sample (63%), over half of the accreted stars within 40 kpc come from a single merger. For most other halos (33%), summing contributions from two mergers is sufficient to reach the 50% threshold, and for a minority (4%) of the halos it takes three mergers to reach this threshold, with no halos requiring four or more mergers. These statistics are summarized in the left panel of Figure 2, which shows the fraction of MWs for which 50% of the stellar halo is comprised of one, two, and three mergers. It is therefore not rare for an MW analog’s halo to be comprised predominantly of a single merger, but it is also not rare for two mergers to together comprise the majority of its ex situ population. In this work, we consider GSE analogs from both of these cases.
@@
Figure 2: (Left:) The number of mergers required to reach the threshold of 50% of ex situ stars within 40 kpc of the galactic center. Approximately 60% of the MW analogs have a single merger that contributes over half of the ex situ stars within 40 kpc, and an additional $\sim 30\%$ reach this threshold with two mergers. It is therefore not unreasonable to expect that an ex situ halo could be predominantly comprised of two mergers, though typically there is a single large merger that comprises the majority of the inner halo. (Right:) The velocity anisotropy $\beta$ of ex situ stellar debris is shown for all large mergers (those that contribute over 50% of their host’s ex situ stellar population within 40 kpc) in red. Those that are considered RA debris lie in the shaded large-$\beta$ region. Each pair of mergers that together comprise a fraction $f>50\%$ is shown in purple. There are occasionally multiple choices of merger pairs in the same halo that satisfy the RA debris criteria (i.e., that fall in the shaded region). There are 21 single-halo RA mergers, and there are 16 RA pairs of mergers selected from 11 MW analogs (five MW analogs have multiple choices of RA merger pairs).
III.2 GSE Analog Selection
@@
Of the mergers identified in TNG, we define a GSE-like subsample of merger debris that
@@
An individual merger that satisfies both criteria is deemed an “RA merger.” In the halos where no single merger yields $f>0.5$ (regardless of its anisotropy), pairs of mergers are considered. In these cases, any two mergers whose stars together satisfy $f>0.5$ and $\beta>0.5$ are deemed an “RA pair.” Stars that are from either an RA merger or an RA pair are referred to as “RA debris.”
These criteria are generally influenced by considerations of the GSE debris in our Galaxy, as the GSE comprises the majority of ex situ material in the inner halo and has a high velocity anisotropy. Criterion 1 ensures that the GSE analog contributes significantly to the inner halo, with no other large contributions, as inferred observationally. Criterion 2 ensures that the debris qualitatively resembles the GSE kinematically, though this criterion is especially permissive.
Of the 98 MW analogs, 21 have a single RA merger and 11 have an RA pair. [^9] We have verified that the RA pairs are not accreted together as part of a larger group structure. The choice of RA pair is not necessarily unique: in five halos, there are two combinations that each satisfy the requirements of an RA pair. Each combination is taken independently, though there is consistently one merger shared between the two pairs. All pairs will be specially marked in the figures, with icons denoting the MW analogs to which they belong.
@@
Notably, individual mergers tend not to have as extreme an anisotropy as the MW stellar halo; the GSE debris in the MW has $\beta\sim 0.8$ (Belokurov et al., 2018; Myeong et al., 2018; Lancaster et al., 2019; Iorio and Belokurov, 2021), which is at the far upper end of what is seen in TNG. Of the debris with $f>0.5$ (i.e., the mergers shown in the right panel of Figure 2), 5% of single mergers and only 2.5% of merger pairs have $\beta>0.8$. This implies the high anisotropy seen in the MW is quite rare.
Considering two mergers together as one set of debris tends to reduce the measured anisotropy, as the mergers do not necessarily have any correlation in velocity space; this is reflected in the shift toward isotropy ($\beta=0$) in the right panel of Figure 2. Despite this, there are still many pairs of mergers for which the contributed stars satisfy both the high-$\beta$ and high-$f$ requirements. This suggests that the presence of such an anisotropic structure cannot immediately preclude the possibility of a multicomponent model. The rest of this work characterizes and aims to distinguish individual GSE-like RA mergers from RA pairs.
@@
III.3 Progenitors of GSE-like Debris
Various properties of GSE analogs are shown in Table 2, for both the RA mergers (top section) and RA pairs (bottom section). The columns include
  1. 1.
    MW ID, the Subfind ID of the MW analog to which the GSE analog belongs;
  2. 2.
    $f$, the fractional contribution of ex situ stars in the inner 40 kpc belonging to the GSE analog;
  3. 3.
@@ III.3 Progenitors of GSE-like Debris
Figure 3: The infall times $t_{\mathrm{infall}}$ and masses $M_{\mathrm{dyn}}$ of all mergers across the TNG MW analogs. The RA mergers passing the selection criteria established in Section III.2 are shown in red, while RA pairs are shown in purple (for the larger-$f$ merger) and pink (for the smaller-$f$ merger), with marker shapes indicating the MW analog to which they belong. Other (non-RA) mergers are shown in black, with the opacity equal to $f$, with higher-$f$ mergers represented by darker markers. The corresponding marginal distributions are shown above and to the right of the main panel. Generally, RA mergers are more recently accreted and larger in mass than those comprising RA pairs. For the RA pairs, the larger-$f$ mergers tend to be more recently accreted and have slightly larger mass.
@@
Figure 4: The distribution of spherical velocities $v_{r}$ and $v_{\phi}$ for RA merger debris in four example MW analogs. The gray-scale histogram shows the full distribution of RA merger debris. The top row shows these distributions for single RA mergers, with 68% of the stars enclosed within the red contour. The bottom row shows the distributions for RA pairs, with 68% of stars from the high-$f$ (low-$f$) merger enclosed within the purple (pink) contour. In each bottom panel, the fractions $f$ are provided in the lower left corner, and an icon unique to the MW analog is provided in the lower right corner, consistent with Figure 3. The examples chosen here highlight the diversity in kinematic features observed in MW’s with RA debris. Versions of this figure for each choice of RA debris are provided in C.
@@
Figure 5: Total energy shown against the $z$-component of angular momentum for the same MW analogs as in Figure 4, with identical formatting, except as noted. Each axis has been normalized to an energy and angular momentum scale characteristic of the RA debris to emphasize the shapes of these distributions. Specifically, $\bar{E}$ is the median energy of RA debris, while $\Delta E$ is the difference between the 2nd and 98th percentile. Similarly, $\Delta L_{z}$ is the difference between the 2nd and 98th percentile values of the RA debris’ $L_{z}$. As in the $v_{r}-v_{\phi}$ plane, these distributions exhibit varying degrees of complexity regardless of whether they correspond to single or pairs of mergers. Versions of this figure for each choice of RA debris are provided in C.
@@
Across the sample, there are only three cases in which an infalling GSE-like galaxy survives to $z=0$. This happens once for a single RA merger (in ID 515296) and twice for the larger-$f$ partner of an RA pair (in MWs 392277 and 394621). In all cases where the GSE-like debris is built up predominantly from a single merger, that event is the most recent major merger (i.e., no subsequent merger has larger $M_{\mathrm{dyn}}/\mathrm{MW}~M_{\mathrm{dyn}}$). However, there are five MWs that accrete an RA pair and subsequently have a merger that is more significant than either in the pair. This happens in IDs 392277, 394621, 446665, 459557, and 535050.
@@
Further elements of the table are summarized in Figure 3, which shows the distribution of infall times, $t_{\mathrm{infall}}$, and total masses at infall, $M_{\mathrm{dyn}}$, for RA mergers (in red), RA pairs (in purple and pink, with the larger-$f$ merger in purple), and all other mergers (in black, with opacity equal to the fraction $f$) across the sample. The marginal distributions of these parameters for the RA debris are shown as normalized histograms on each axis.